# An Improved Machine Learning Approach for RFI Mitigation in FAST-SETI Survey Archival Data

### v1

@@ 1 Introduction
Whether the Earth is the only host of life in the universe has long been a question captivating humanity. Relying on the Copernican principle and the Drake equation (drake1961project), the majority of scientists hold the view that intelligent life must exist beyond Earth (namely Extraterrestrial Intelligence, or ETI for short; 2025NatAs...9...16V). There are three primary approaches to search for extraterrestrial life: (1) direct in situ detection of biosignatures in specific environments (2015Sci...347..415W); (2) remote sensing of biological signals from the atmospheres and surfaces of exoplanets (2014Sci...343..171R; 2014PNAS..11112634S); and (3) detect the technosignatures via the search for extraterrestrial intelligent, i.e. SETI (1959Natur.184..844C). From an astronomical viewpoint, the detection of biosignatures faces enormous challenges due to the constraints in detection range and observation time (2005AsBio...5..706S; 2014ApJ...781...54R; 2016ApJ...819L..13S). In contrast, SETI is only required to detect technological signals that are either intentionally or unintentionally emitted by an ETI. Most SETI researches are typically conducted in the radio wave band because radio signals propagate effectively through interstellar space. With advancements in radio instrumentation, the available bandwidth of the radio SETI systems has expanded to tens of GHz in recent years (2018PASP..130d4502M).
@@
Due to broadening effects on electromagnetic signals in nature, the narrowest radio spectrum of natural astrophysical phenomena has a width of at least approximately 500 Hz (1987MNRAS.225..491C). Therefore, narrowband signals ($\sim$ Hz) are highly suitable for ETI to carry information (2001ARA&A..39..511T). When a narrowband signal is transmitted from a distant source, the relative motion between the transmitter and receiver causes a frequency shift known as Doppler drift. Therefore, the narrowband drifting signal is considered a target in SETI research.
@@
The SETI radio observations predominantly rely on two ways: commensal sky surveys and targeted observations (2001ARA&A..39..511T). The targeted observations focus on pre-determined objects, most commonly nearby stars, while the commensal sky surveys scan large areas of sky to find potential ETI signals for further examination. For single-dish telescopes, commensal sky surveys typically employ the “drifting scan”observation strategy, leveraging the Earth’s rotation to perform systematic scans across the right ascension (R.A.). The SETI commensal surveys are represented by projects such as SERENDIP program, SETI@home (2001SPIE.4273..104W), the Five-hundred-meter Aperture Spherical radio Telescope (FAST) SETI backend (zhang2020first; wang2023search), and COSMIC (2024AJ....167...35T). In recent years, targeted SETI observations are increasingly being conducted (e.g., 2013ApJ...767...94S; 2017ApJ...849..104E; 2017AJ....153..110G; 2016ApJ...827L..22T; 2018ApJ...856...31T; 2016AJ....152..181H; 2018ApJ...869...66H; 2020AJ....160..162H; 2019AJ....157..122P; 2020AJ....159...86P; 2020AJ....160...29S; 2020PASA...37...35T; 2021NatAs...5.1148S; 2021AJ....161..286T; Gajjar_2021; tao2022sensitive; 2023AJ....166..190T; 2023AJ....166..245H; 2023AJ....165..132L; Luan_2025). Nevertheless, SETI commensal surveys can still serve as a complement to targeted observations, as they offer a few advantages: (1) they scan larger sky regions for ETI signals; (2) their observation durations are orders of magnitude longer; and (3) they are target-agnostic, thereby mitigating anthropocentric biases in target selection.
@@
As the largest single-aperture radio telescope on Earth, FAST (2000ASPC..213..523N; 2011IJMPD..20..989N; 2016RaSc...51.1060L; 2019SCPMA..6259502J; 2020RAA....20...64J) provides us with great opportunities to SETI observations (nan2006five; 2020RAA....20...78L). With its L-band (1.051.451.05-1.45 GHz) 19-beam receiver, FAST can cover a large sky area (declinations from $14^{\circ}.34$ to $65^{\circ}.7$) and exhibits extremely high sensitivity ($\sim 2000\;m^{2}\;K^{-1}$). The first FAST SETI commensal survey, a drift-scan observation conducted during its commissioning in July 2019, was analyzed by zhang2020first, who identified two groups of high-confidence ETI candidate signals, and later by wang2023search, who identified 14 groups of ETI candidate signals from the same dataset. FAST has also conducted multiple targeted SETI observations in recent years (tao2022sensitive; 2023AJ....166..190T; 2023AJ....166..245H; 2023AJ....165..132L; Luan_2025). In the future, FAST will conduct more SETI observations, comprising both targeted searches and commensal surveys.
@@
The major challenge of radio SETI observations is identifying and mitigating radio frequency interference (RFI) caused by human technology. In previous analyses of FAST-SETI commensal survey data, zhang2020first and wang2023search employed the Nebula platform  [^1] http://setiathome.berkeley.edu/nebula and Hough transform method, respectively, to mitigate substantial RFI. To further remove residual RFI, both studies integrated a machine learning approach, i.e. the K-Nearest Neighbor (KNN) algorithm. Although both studies mentioned above have successfully mitigated most RFI, it is still meaningful to further improve the RFI mitigation algorithms in the following two aspects: (1) removing more RFI to reduce the work of visual inspection; (2) improving computational efficiency to accelerate data processing.
@@
In this paper, we propose an improved unsupervised machine learning approach—the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (ester1996density; schubert2017dbscan)—to replace the KNN algorithm for mitigating residual RFI. Moreover, we present the results of applying the DBSCAN algorithm to the same FAST-SETI commensal survey data analyzed in zhang2020first and wang2023search. In Section 2, we describe the RFI removal methods proposed and used in this paper in detail. Then we briefly introduce the data and present the results of RFI removal, candidate selection and analysis in Section 3. We finally conclude this work in Section 5.
@@
Figure 1: The waterfall (frequency-time) plot of the raw observational data (i.e., hits, marked with black dots) from the first 1800s of observation. The narrowband RFI is prominent, appearing as vertical lines; some broadband RFI (appearing as horizontal line segments) also exists. This figure is same as the Figure 1 in wang2023search.
2 Methods for RFI Mitigation
@@
In the SETI commensal survey, RFI removal algorithms are primarily applied to process “hit” records as the input data. A “hit” in this paper refers to information contained in potential signal of interest that has a high signal-to-noise ratio (S/N) in its frequency channels at each moment (See Section 3.1 and zhang2020first for details). Each hit contains multiple observational parameters, including timestamp (UTC), frequency channel identifier, telescope pointing coordinates, and system temperature measurements, which can be used to identify RFI. As shown in Figure 1, all hits are distributed on a waterfall plot of time $t$ and frequency $f$. Most hits are RFI, while some may be interesting ETI candidates. Unlike targeted observations that use “filterbank” data with complete spectral information (tao2022sensitive; 2023AJ....166..190T; 2023AJ....166..245H; 2023AJ....165..132L; Luan_2025), commensal surveys typically involve long-duration observations containing only hit data. Therefore, specialized software for analyzing hit data is required.
@@
2.1 Persistent and Drifting Narrowband RFI Removal
@@
We remove the first two types of RFI following the methods proposed in wang2023search. Persistent narrowband RFI typically manifests as signals within a specific frequency channel that appear across a wide expanse of the sky or endure for extended durations. Such signals are generally not of astronomical origin and appear as prominent vertical lines in Figure 1. In the SETI commensal survey observations, this persistence often translates to detection over a broad sky area as the Earth’s rotation moves the telescope’s pointing. For persistent narrowband RFI removal, the full frequency range (1000–1500 MHz) is divided into small bins (each $\sim$ 7.45 Hz in size). A sky angular separation threshold is then established as 1.5 times the distance between adjacent beam centers. When any pair of hits in a frequency bin exhibits a sky angular separation exceeding this threshold, the signal in entire frequency bin is marked as RFI (see Section 2.1 of wang2023search for details). The characteristic of drifting (narrowband) RFI is that drifts in frequency. Such signals, which can originate from sources like satellites, moving objects, or local oscillator malfunctions, often appear as slanted or curved lines in Figure 1. For drifting (narrowband) RFI removal, the frequency-time plane is divided into a frequency window of 20 MHz and a time window of 600 seconds. Within each such window, the hits converted to binary images with pixel sizes of 0.004 MHz and 20 seconds. Line segment detection is then performed on these binary images by the OpenCV13OpenCV13222https://opencv.org/ implementation of the probabilistic Hough transform, $HoughLinesP$, which is based on the progressive probabilistic Hough transform (PPHT) proposed by matas2000robust, and any detected line segments are extended by several pixels at each end, and all hits located within a “corridor” of less than several pixels from these extended lines are identified and removed as drifting RFI (see Section 2.2 of wang2023search for details).
2.2 The DBSCAN Algorithm for Residual RFI Removal
@@
After removing the persistent and drifting narrowband RFI as described above, there is still a small amount of residual RFI in the data. Residual RFI typically has multiple types, and two typical examples are narrowband RFI and broadband RFI. For narrowband RFI, we cannot completely remove it using the method of Section 2.1 due to the fact that its power is sometimes below our threshold. For broadband RFI, it is easy to remove if the bandwidth is very large. However, if the bandwidth is less than several MHz, it is extremely difficult to detect using the traditional methods.
@@
In the SETI commensal survey, the ETI signals have two limitations: they cannot last a long time (because of the pointing direction drift of the telescope, as discussed in Section 3.4) or cover a wide frequency range (because of the commonly assumed narrowband nature). Thus, ETI signals form smaller clusters than RFI in time and frequency. So, the clustering algorithm can be used to remove residual RFI. In previous studies, zhang2020first and wang2023search employed the KNN algorithm to find the nearest 100 hits for each hit and calculate the mean distance. However, the KNN algorithm relies on a global density threshold to classify hits, which makes it insensitive to locally sparse RFI and challenging to identify irregularly shaped hit clusters. Moreover, the computational complexity of the KNN algorithm is O (${n}^{2}$), which becomes particularly time-consuming when dealing with massive data, potentially failing to meet real-time processing demands.
@@
We apply an improved machine learning method, DBSCAN algorithm (ester1996density), to remove the residual RFI. DBSCAN is an unsupervised learning method, meaning it identifies inherent patterns directly from unlabeled data, which is ideal for SETI research where verified ETI labels are unavailable. This differs from the previously used KNN algorithm, which is conventionally a supervised learning algorithm but is applied for an unsupervised task by applying a threshold to the mean k-nearest neighbor distance of each hit to identify and remove large, dense clusters considered RFI (wang2023search; zhang2020first). The primary advantage of DBSCAN is using local density to define clusters, the derived unsupervised algorithm (HDBCSAN algorithm) has been successfully applied to RFI classification in other recent technosignature searches (jacobson2025anomaly). DBSCAN is effective in separating interference signal clusters in the noise space while preserving sparsely distributed potential ETI signals by identifying arbitrarily shaped clusters in high-density regions. Due to the hits are irregular in time and frequency scales, making the clusters formed by RFI also irregular in shape. The DBSCAN algorithm is particularly well-suited to detect these irregular clusters. With the help of the robustness of the DBSCAN algorithm in noisy data, the data noise points can be effectively identified, that is, the signals that do not belong to RFI can be found.
@@
    The hit2 lies within the Eps-neighborhood(as defined by the ester1996density) of hit1.
@@
3 Data and results
3.1 Data and Preprocessing
@@
To verify the effectiveness of the RFI removal method, simulated ETI signals, called “birdies”, were injected into the observational data. These simulated signals comprise 20 groups with a total of 294 signals, generated and described in zhang2020first and wang2023search. Therefore, 47779 hits (including birdies) are obtained to verify the proposed method. The simulated signals were generated under controlled conditions emulating celestial point sources with single-frequency channel emissions, strategically positioned along FAST’s observational trajectory.
@@
3.2 RFI Removal Results
According to Section 2.1, we first remove the persistent and drifting narrowband RFI, which account for approximately 99.9912% of all data. After this removal, a total of 47779 residual hits remain, including 294 birdies. The Figure 2 shows the results after removing these two types of RFI.
For the residual RFI removal, we only focus on the two features of hits, namely time and frequency. Given the very wide ranges of both time and frequency, we first apply linear normalization to standardize time and frequency into the range [0,1] [^3] The 5-hour observation duration falls within the typical observation range and excludes extreme cases in the FAST-SETI commensal surveys. For these observations, we applied the [0,1] normalization to standardize time and frequency values, making the data easier to process and compare. This ensures consistent results for different durations within a reasonable range, as minor changes in observation duration would not significantly affect the analysis when using this normalization method.. Then, we adjust the parameters Eps and MinPts in the DBSCAN algorithm. To identify the large clusters formed by the residual RFI, the value of MinPts is set to 55444We adjusted MinPts from 35 to 95 and found that MinPts = 55 resulted in the highest residual RFI removal rate.. The selection of Eps is determined by the simulated signal “birdies”, as shown in Figure 3, when the Eps is set to ${7.5}\times{10}^{-3}$, the loss rate of the “birdies” is 0, and the number of hits to be removed is the largest. Based on the parameters described above, we have successfully identified and removed 36977 residual RFI (77.87% of the residual hits) using the DBSCAN algorithm (see the left panel in Figure 4). This performance is significantly higher than that of the KNN algorithm, which removed 33445 residual RFI (70.43% of the residual hits) under identical processing conditions and the same dataset (see the right panel in Figure 4, wang2023search). After the residual RFI removal, the number of valid hits is 10508. Data processing using the DBSCAN algorithm is implemented with the scikit-learn package in Python (pedregosa2011scikit).

### v2

@@ I Introduction
Whether the Earth is the only host of life in the universe has long been a question captivating humanity. Relying on the Copernican principle and the Drake equation (Drake, 1961), the majority of scientists hold the view that intelligent life must exist beyond Earth (namely Extraterrestrial Intelligence, or ETI for short; Vickers et al. (2025)). There are three primary approaches to search for extraterrestrial life: (1) direct in situ detection of biosignatures in specific environments (Webster et al., 2015); (2) remote sensing of biological signals from the atmospheres and surfaces of exoplanets (Roth et al., 2014; Seager, 2014); and (3) detect the technosignatures via the search for extraterrestrial intelligent, i.e. SETI (Cocconi and Morrison, 1959). From an astronomical viewpoint, the detection of biosignatures faces enormous challenges due to the constraints in detection range and observation time (Segura et al., 2005; Rodler and López-Morales, 2014; Schwieterman et al., 2016). In contrast, SETI is only required to detect technological signals that are either intentionally or unintentionally emitted by an ETI. Most SETI researches are typically conducted in the radio wave band because radio signals propagate effectively through interstellar space. With advancements in radio instrumentation, the available bandwidth of the radio SETI systems has expanded to tens of GHz in recent years (MacMahon et al., 2018).
@@
Due to broadening effects on electromagnetic signals in nature, the narrowest radio spectrum of natural astrophysical phenomena has a width of at least approximately 500 Hz (Cohen et al., 1987). Therefore, narrowband signals ($\sim$ Hz) are highly suitable for ETI to carry information (Tarter, 2001). When a narrowband signal is transmitted from a distant source, the relative motion between the transmitter and receiver causes a frequency shift known as Doppler drift. Therefore, the narrowband drifting signal is considered a target in SETI research.
@@
The SETI radio observations predominantly rely on two ways: commensal sky surveys and targeted observations (Tarter, 2001). The targeted observations focus on pre-determined objects, most commonly nearby stars, while the commensal sky surveys scan large areas of sky to find potential ETI signals for further examination. For single-dish telescopes, commensal sky surveys typically employ the “drifting scan”observation strategy, leveraging the Earth’s rotation to perform systematic scans across the right ascension (R.A.). The SETI commensal surveys are represented by projects such as SERENDIP program, SETI@home (Werthimer et al., 2001), the Five-hundred-meter Aperture Spherical radio Telescope (FAST) SETI backend (Zhang et al., 2020; Wang et al., 2023), and COSMIC (Tremblay et al., 2024). In recent years, targeted SETI observations are increasingly being conducted (e.g., Siemion et al. (2013); Enriquez et al. (2017); Gray and Mooley (2017); Tingay et al. (2016, 2018); Harp et al. (2016, 2018, 2020); Pinchuk et al. (2019); Price et al. (2020); Sheikh et al. (2020); Tremblay and Tingay (2020); Smith et al. (2021); Traas et al. (2021); Gajjar et al. (2021); Tao et al. (2022, 2023); Huang et al. (2023); Luan et al. (2023, 2025)). Nevertheless, SETI commensal surveys can still serve as a complement to targeted observations, as they offer a few advantages: (1) they scan larger sky regions for ETI signals; (2) their observation durations are orders of magnitude longer; and (3) they are target-agnostic, thereby mitigating anthropocentric biases in target selection.
@@
As the largest single-aperture radio telescope on Earth, FAST (Nan et al., 2000, 2011; Li and Pan, 2016; Jiang et al., 2019, 2020) provides us with great opportunities to SETI observations (Nan, 2006; Li et al., 2020). With its L-band (1.051.451.05-1.45 GHz) 19-beam receiver, FAST can cover a large sky area (declinations from $14^{\circ}.34$ to $65^{\circ}.7$) and exhibits extremely high sensitivity ($\sim 2000\;m^{2}\;K^{-1}$). The first FAST SETI commensal survey, a drift-scan observation conducted during its commissioning in July 2019, was analyzed by Zhang et al. (2020), who identified two groups of high-confidence ETI candidate signals, and later by Wang et al. (2023), who identified 14 groups of ETI candidate signals from the same dataset. FAST has also conducted multiple targeted SETI observations in recent years (Tao et al., 2022, 2023; Huang et al., 2023; Luan et al., 2023, 2025). In the future, FAST will conduct more SETI observations, comprising both targeted searches and commensal surveys.
@@
The major challenge of radio SETI observations is identifying and mitigating radio frequency interference (RFI) caused by human technology. In previous analyses of FAST-SETI commensal survey data, Zhang et al. (2020) and Wang et al. (2023) employed the Nebula platform  [^1] http://setiathome.berkeley.edu/nebula and Hough transform method, respectively, to mitigate substantial RFI. To further remove residual RFI, both studies integrated a machine learning approach, i.e. the K-Nearest Neighbor (KNN) algorithm. Although both studies mentioned above have successfully mitigated most RFI, it is still meaningful to further improve the RFI mitigation algorithms in the following two aspects: (1) removing more RFI to reduce the work of visual inspection; (2) improving computational efficiency to accelerate data processing.
@@
In this paper, we propose an improved unsupervised machine learning approach—the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996; Schubert et al., 2017)—to replace the KNN algorithm for mitigating residual RFI. Moreover, we present the results of applying the DBSCAN algorithm to the same FAST-SETI commensal survey data analyzed in Zhang et al. (2020) and Wang et al. (2023). In Section II, we describe the RFI removal methods proposed and used in this paper in detail. Then we briefly introduce the data and present the results of RFI removal, candidate selection and analysis in Section III. We finally conclude this work in Section V.
@@
Figure 1: The waterfall (frequency-time) plot of the raw observational data (i.e., hits, marked with black dots) from the first 1800s of observation. The narrowband RFI is prominent, appearing as vertical lines; some broadband RFI (appearing as horizontal line segments) also exists. This figure is same as the Figure 1 in Wang et al. (2023).
II Methods for RFI Mitigation
@@
In the SETI commensal survey, RFI removal algorithms are primarily applied to process “hit” records as the input data. A “hit” in this paper refers to information contained in potential signal of interest that has a high signal-to-noise ratio (S/N) in its frequency channels at each moment (See Section III.1 and Zhang et al. (2020) for details). Each hit contains multiple observational parameters, including timestamp (UTC), frequency channel identifier, telescope pointing coordinates, and system temperature measurements, which can be used to identify RFI. As shown in Figure 1, all hits are distributed on a waterfall plot of time $t$ and frequency $f$. Most hits are RFI, while some may be interesting ETI candidates. Unlike targeted observations that use “filterbank” data with complete spectral information (Tao et al., 2022, 2023; Huang et al., 2023; Luan et al., 2023, 2025), commensal surveys typically involve long-duration observations containing only hit data. Therefore, specialized software for analyzing hit data is required.
@@
II.1 Persistent and Drifting Narrowband RFI Removal
@@
We remove the first two types of RFI following the methods proposed in Wang et al. (2023). Persistent narrowband RFI typically manifests as signals within a specific frequency channel that appear across a wide expanse of the sky or endure for extended durations. Such signals are generally not of astronomical origin and appear as prominent vertical lines in Figure 1. In the SETI commensal survey observations, this persistence often translates to detection over a broad sky area as the Earth’s rotation moves the telescope’s pointing. For persistent narrowband RFI removal, the full frequency range (1000–1500 MHz) is divided into small bins (each $\sim$ 7.45 Hz in size). A sky angular separation threshold is then established as 1.5 times the distance between adjacent beam centers. When any pair of hits in a frequency bin exhibits a sky angular separation exceeding this threshold, the signal in entire frequency bin is marked as RFI (see Section 2.1 of Wang et al. (2023) for details). The characteristic of drifting (narrowband) RFI is that drifts in frequency. Such signals, which can originate from sources like satellites, moving objects, or local oscillator malfunctions, often appear as slanted or curved lines in Figure 1. For drifting (narrowband) RFI removal, the frequency-time plane is divided into a frequency window of 20 MHz and a time window of 600 seconds. Within each such window, the hits converted to binary images with pixel sizes of 0.004 MHz and 20 seconds. Line segment detection is then performed on these binary images by the OpenCV13OpenCV13222https://opencv.org/ implementation of the probabilistic Hough transform, $HoughLinesP$, which is based on the progressive probabilistic Hough transform (PPHT) proposed by Matas et al. (2000), and any detected line segments are extended by several pixels at each end, and all hits located within a “corridor” of less than several pixels from these extended lines are identified and removed as drifting RFI (see Section 2.2 of Wang et al. (2023) for details).
II.2 The DBSCAN Algorithm for Residual RFI Removal
@@
After removing the persistent and drifting narrowband RFI as described above, there is still a small amount of residual RFI in the data. Residual RFI typically has multiple types, and two typical examples are narrowband RFI and broadband RFI. For narrowband RFI, we cannot completely remove it using the method of Section II.1 due to the fact that its power is sometimes below our threshold. For broadband RFI, it is easy to remove if the bandwidth is very large. However, if the bandwidth is less than several MHz, it is extremely difficult to detect using the traditional methods.
@@
In the SETI commensal survey, the ETI signals have two limitations: they cannot last a long time (because of the pointing direction drift of the telescope, as discussed in Section III.4) or cover a wide frequency range (because of the commonly assumed narrowband nature). Thus, ETI signals form smaller clusters than RFI in time and frequency. So, the clustering algorithm can be used to remove residual RFI. In previous studies, Zhang et al. (2020) and Wang et al. (2023) employed the KNN algorithm to find the nearest 100 hits for each hit and calculate the mean distance. However, the KNN algorithm relies on a global density threshold to classify hits, which makes it insensitive to locally sparse RFI and challenging to identify irregularly shaped hit clusters. Moreover, the computational complexity of the KNN algorithm is O (${n}^{2}$), which becomes particularly time-consuming when dealing with massive data, potentially failing to meet real-time processing demands.
@@
We apply an improved machine learning method, DBSCAN algorithm (Ester et al., 1996), to remove the residual RFI. DBSCAN is an unsupervised learning method, meaning it identifies inherent patterns directly from unlabeled data, which is ideal for SETI research where verified ETI labels are unavailable. This differs from the previously used KNN algorithm, which is conventionally a supervised learning algorithm but is applied for an unsupervised task by applying a threshold to the mean k-nearest neighbor distance of each hit to identify and remove large, dense clusters considered RFI (Wang et al., 2023; Zhang et al., 2020). The primary advantage of DBSCAN is using local density to define clusters, the derived unsupervised algorithm (HDBCSAN algorithm) has been successfully applied to RFI classification in other recent technosignature searches (Jacobson-Bell et al., 2025). DBSCAN is effective in separating interference signal clusters in the noise space while preserving sparsely distributed potential ETI signals by identifying arbitrarily shaped clusters in high-density regions. Due to the hits are irregular in time and frequency scales, making the clusters formed by RFI also irregular in shape. The DBSCAN algorithm is particularly well-suited to detect these irregular clusters. With the help of the robustness of the DBSCAN algorithm in noisy data, the data noise points can be effectively identified, that is, the signals that do not belong to RFI can be found.
@@
    The hit2 lies within the Eps-neighborhood(as defined by the Ester et al. (1996)) of hit1.
@@
III Data and results
III.1 Data and Preprocessing
@@
To verify the effectiveness of the RFI removal method, simulated ETI signals, called “birdies”, were injected into the observational data. These simulated signals comprise 20 groups with a total of 294 signals, generated and described in Zhang et al. (2020) and Wang et al. (2023). Therefore, 47779 hits (including birdies) are obtained to verify the proposed method. The simulated signals were generated under controlled conditions emulating celestial point sources with single-frequency channel emissions, strategically positioned along FAST’s observational trajectory.
@@
III.2 RFI Removal Results
According to Section II.1, we first remove the persistent and drifting narrowband RFI, which account for approximately 99.9912% of all data. After this removal, a total of 47779 residual hits remain, including 294 birdies. The Figure 2 shows the results after removing these two types of RFI.
For the residual RFI removal, we only focus on the two features of hits, namely time and frequency. Given the very wide ranges of both time and frequency, we first apply linear normalization to standardize time and frequency into the range [0,1] [^3] The 5-hour observation duration falls within the typical observation range and excludes extreme cases in the FAST-SETI commensal surveys. For these observations, we applied the [0,1] normalization to standardize time and frequency values, making the data easier to process and compare. This ensures consistent results for different durations within a reasonable range, as minor changes in observation duration would not significantly affect the analysis when using this normalization method.. Then, we adjust the parameters Eps and MinPts in the DBSCAN algorithm. To identify the large clusters formed by the residual RFI, the value of MinPts is set to 55444We adjusted MinPts from 35 to 95 and found that MinPts = 55 resulted in the highest residual RFI removal rate.. The selection of Eps is determined by the simulated signal “birdies”, as shown in Figure 3, when the Eps is set to ${7.5}\times{10}^{-3}$, the loss rate of the “birdies” is 0, and the number of hits to be removed is the largest. Based on the parameters described above, we have successfully identified and removed 36977 residual RFI (77.87% of the residual hits) using the DBSCAN algorithm (see the left panel in Figure 4). This performance is significantly higher than that of the KNN algorithm, which removed 33445 residual RFI (70.43% of the residual hits) under identical processing conditions and the same dataset (see the right panel in Figure 4, Wang et al. (2023)). After the residual RFI removal, the number of valid hits is 10508. Data processing using the DBSCAN algorithm is implemented with the scikit-learn package in Python (Pedregosa et al., 2011).